An observer at rest near the source of the sound with f = 784 Hz. Another source of sound with 775 Hz is moving toward the observer at 2 m/s. If the v of the sound waves in air is 340 m/s, what is the beat f heard by the observer?

1
Expert's answer
2022-02-02T14:38:00-0500

The frequency of the second source heard by the observer:


"f=775\dfrac{340}{340-2}=779.6\text{ Hz}."

Beat frequency:


"f_b=|779.6-784|=4.4\text{ Hz}."